OpenStudy (anonymous):

I WILL MEDAL AND FAN! :) Which is the direct linear equation for the relationship? Y varies directly with x and y=12 when x=3 a: y=4x b: y=1/4 x c: y= x+9 d: y= x-9

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@dan815

OpenStudy (freckles):

y varies directly with x looks like y=kx find k using (x,y)=(3,12)

OpenStudy (anonymous):

k=6

OpenStudy (freckles):

so you say saying 12=6(3) that is impossibly

OpenStudy (anonymous):

no sorry I was dividing

OpenStudy (freckles):

well you should be able to get the correct k by dividing you have 12=k*3 how do you solve for k

OpenStudy (freckles):

you said 6 which is impossible because 3(6)=18 not 12

OpenStudy (anonymous):

so its 4?

OpenStudy (anonymous):

so that makes it A correct?

OpenStudy (anonymous):

it would be A or C

OpenStudy (freckles):

it would be A since y=kx is a direct variation and we want to find k given y=12 and x=3 12=3k dividing 3 on both sides 12/3=k 4=k so that means instead of y=kx you write y=4x since you found k to be 4